The position of a squirel running in a park s gven The position of a squirrel running in a park is given by (0.280 m/s)t + (0.0360 m/t) + (0.0190 m/s3)tj

A)

B)

C) At 5.24 s , how far is the squirrel from its initial position?

D) At 5.24 s , what is the magnitude of the squirrel's velocity?

E) At 5.24 s , what is the direction (in degrees counterclockwise from +x-axis) of the squirrel's velocity?

Express your answer to three significant figures and include the appropriate units.
Answer #1

given


position vector = r = (0.28t + 0.036 t^2 ) i + 0.019 t^3 j

velocity = v = dr/dt = (0.28 + 2*0.036 t ) i + (3*0.019 t^2) j


A)

Vx = (0.28 + 0.072 t )  

B) vy = + (0.057 t^2 )

c)

at t= 5.24 s

r = ((0.28*5.24) + (0.036*5.24^2) ) i + (0.019*5.24^3) j

r = 2.46 i + 2.73 j


distane = r = 3.67 m


d)


vx = 0.656 m/s

vy = 1.56 m/s

magnitude = sqrt(vx^2+vy^2) = sqrt(0.656^2+1.56^2) = 1.69 m


(E)

direction = tan^-1(vy/vx) = 67.2 degrees counterclockwise from +x-axis